# InstructRAG: Instructing Retrieval-Augmented Generation with Explicit Denoising

### v1

@@ 1 Introduction
In light of this, retrieval-augmented generation (RAG) [3, 18, 24, 35, 45] has been introduced to enhance the generation accuracy of LMs in knowledge-intensive tasks by leveraging the most up-to-date information and specialized knowledge from external sources [34, 74, 87, 100]. However, the retrieved contents are typically mixed with irrelevant or even erroneous information due to the absence of perfect retrieval solutions [22, 33, 36, 37, 65] and the presence of noisy data in the retrieval corpus [23, 46], posing a long-standing challenge to almost all RAG systems. Typically, vanilla RAG approaches address this issue implicitly by training LMs to directly predict correct answers despite potentially noisy inputs. Such latent processes are not only difficult to interpret and verify but also vulnerable to higher noise ratios, especially when the number of retrieved documents is large [9, 10, 49, 79]. On the other hand, obtaining high-quality explicit denoising supervision often requires substantial human efforts, which is time-consuming and costly.
@@ 2 Our Method: InstructRAG
During the entire process, InstructRAG does not require any additional supervisions beyond standard RAG methods. By default, we instantiate both and with the same off-the-shelf instruction-tuned model (i.e., meta-llama/Meta-Llama-3-8B-Instruct), making InstructRAG a self-synthesis method. We also experiment with different instantiations of and and conduct ablation study in both training-free and trainable settings (§ 3.3). For simplicity, we use placeholders to represent omitted instructions in the prompts presented in this section, while the full list of complete prompt templates is provided in Appendix B.
@@ 2.1 Problem Setting
We adopt the standard RAG setting where the LM has access to annotated datasets of downstream tasks (e.g., question-answering task ), and an external knowledge base with off-the-shelf retrievers for retrieval. Different from Self-RAG [3] which leverages additional supervision sources like GPT-4 [1], we assume the model has strictly limited access to the above two information sources. Given a question , the retriever returns a set of potentially noisy documents from the external knowledge base. The model is then tasked to predict the correct answer to the given question based on and its own parametric knowledge, denoted as .
@@ 2.2 Rationale Generation via Instruction-Following
Recent studies [44, 56, 59] have made encouraging progress in aligning LMs with human preferences and intentions, enabling the synthesis of high-quality data that closely follows user instructions [84]. Inspired by these advances, we propose to leverage the LM’s strong instruction-following ability to generate explicit denoising responses (i.e., rationales) for RAG. As shown in Table 1, given a QA pair and a set of retrieved documents , we prompt a LM (as the rationale generator) with denoising instructions to produce the corresponding rationale that distinguishes useful documents from noisy ones and explains how the contexts lead to the ground-truth answer . To ensure that the synthetic rationales are aligned with the ground-truth answers, we use a simple substring match to assess their consistency. The consistency ratio on training samples with at least one relevant document containing the ground-truth answer is 98% on average across five benchmarks, supporting the reliability of synthetic rationales as a sanity check. This allows us to effectively augment the standard dataset with self-synthesized denoising rationales solely by instructing the LM, without any additional supervision.
We also validate the necessity of using a generator (i.e., ) to create the rationales instead of employing simple heuristics — without the generator, rationales can be created in a template-based manner (Table 7), by roughly identifying relevant retrieved documents through simple substring-matching with the ground-truth answer. However, as demonstrated in our ablation study, this approach suffers from semantically inaccurate matching of relevant documents, leading to significant performance degradation. Another advantage of the generator is that it can produce high-quality rationales even without referring to the ground-truth answer, which only results in a minor performance drop. More detailed analyses on rationale generation design can be found in our ablation study (§ 3.3).
@@ 3.1 Experimental Setting
RAG tasks and evaluation metrics. We extensively validate the effectiveness of InstructRAG on five knowledge-intensive benchmarks, including PopQA [54], TriviaQA [32], Natural Questions [40], ASQA [69], and 2WikiMultiHopQA [19]. We use Wikipedia corpus as the external retrieval source, and test our method with a wide range of both sparse and dense off-the-shelf retrievers, including BM25 [62], DPR [33], GTR [58] and Contriver [22]. The retrieval quality is measured by Recall@, indicting whether the retrieved documents provide information about the correct answer to the given question. The detailed retrieval configuration and statistics for each dataset are summarized in Table 4. Following standard evaluation settings [3], we adopt the official metric of correctness (str-em), citation precision (pre) and recall (rec) for ASQA [14], and use accuracy for the other tasks, which measures whether the ground-truth answers are included in the model generations [54, 64].
Baselines. We compare our method with a wide range of RAG baselines under both training-free and trainable settings. Given that state-of-the-art LMs have incorporated a large amount of world-knowledge during the pre-training stage, we also report the performance of a non-retrieval baseline (namely, vanilla zero-shot prompting) for reference.
Specifically, the training-free RAG baselines includes: (1) in-context retrieval-augmented language modeling (RALM) [61], a prompting method that directly extends the non-retrieval baseline by presenting the model with retrieved documents; (2) few-shot demonstration with instruction, an ICL method using ground-truth question-answer pairs sampled from the training set as demonstration exemplars.
The trainable RAG baselines include: (1) vanilla supervised fine-tuning (SFT), a supervised learning method with the training objective of maximizing the data likelihood of ground-truth answer; (2) Self-RAG [3], a strong trainable baseline, focusing on adaptive retrieval controlled by special reflection tokens. This method was originally built on Llama-2 [73] with additional supervisions from GPT-4. For a fair comparison, we re-implement Self-RAG on Llama-2 and Llama-2 with augmented training set released by its authors, and report its performance as the higher one between the original scores and our reproduced results. As our method adopts instruction-tuned Llama-3 as the backbone model, we also implement Self-RAG with Llama-3-Instruct and optimize its performance through extensive hyper-parameters search. More implementation details are available in Appendix A, and the inference prompts for each method are presented in Appendix B.
@@ 3.2 Main Result
RAG without training. The second block shows the comparison among training-free RAG methods. It is evident that in-context RALM and few-shot demonstration with instruction methods generally achieve higher performance than the non-retrieval baseline, highlighting the importance of retrieval for knowledge-intensive tasks. Encouragingly, our InstructRAG-ICL consistently outperforms all training-free baselines across various metrics, confirming the effectiveness of self-synthesized denoising rationales. Moreover, the boost from 8B to 70B model also demonstrate that InstructRAG-ICL scales effectively with larger base models, validating the generalizability of our method.
RAG with training. As present in the bottom block of Table 5, our InstructRAG-FT not only significantly outperforms trainable RAG baselines on almost every metric, but also surpasses all training-free and non-retrieval baselines across all five benchmarks. The only exception is in the ASQA task, where our method slightly underperforms Self-RAG in terms of citation metrics (i.e., pre and rec). Note that our work primarily focuses on explicit denoising for RAG to improve the correctness of generations, which is measured by em. Despite not being explicitly optimized for citation metrics, our method still achieves competitive citation performance, significantly enhancing both generation accuracy and trustworthiness. Another interesting finding is that Self-RAG consistently exhibits inferior performance compared to vanilla SFT, and even underperforms the training-free in-context RALM baseline across all benchmarks. We speculate the reason might be that these RAG tasks favor more domain-specific knowledge than general knowledge. However, it is challenging for Self-RAG to directly leverage in-domain features from existing training data as it requires GPT-4 to generate reflection tokens on these benchmarks, which is not available in our problem (§ 2.1).
3.3 Ablation Study
@@ 3.4 Analysis
3.5 Case Study
@@ 4.1 Retrieval-augmented Generation
Retrieval-augmented generation (RAG) is a widely adopted approach to complement the parametric knowledge of large language models (LLMs) with external information sources [2, 4, 5, 15, 18, 35, 45, 61, 65], demonstrating promising potential to reduce hallucinations and enhance the generation accuracy of LLMs across various real-world applications [8, 31, 48, 51, 68, 71, 100]. Recently, a growing research effort has been devoted to enhancing RAG from various aspects, such as improving decoding efficiency [57, 30, 50], exploring long-context retrieval [83, 89], and addressing practical concerns such as adversarial retrieval [80, 99, 101] and privacy leakage [21, 92]. Despite their advantages, these RAG systems inevitably suffer from irrelevant information introduced by imperfect retrievers or noisy retrieval corpora. However, most existing works typically address this issue by improving the retrieval quality and reducing noise exposure to the model [17, 27, 63, 76, 86, 87, 93, 95]. Notable methods include adaptive retrieval [3, 29, 88] and query rewriting [7, 55]. In contrast, our work focuses on an orthogonal direction that aims to develop explicit denoising methods for RAG, thereby enhancing the model’s noise robustness and generation accuracy, even in highly noisy contexts.
@@ 4.2 Eliciting Reasoning in Large Language Models
Recent studies have extensively explored the reasoning capability of LMs, but typically not in the context of RAG where potentially noisy retrieved contents may mislead the reasoning if not properly addressed. Chain-of-thought (CoT) prompting [78] is a simple yet effective method to elicit complex, step-by-step reasoning from LMs by showing exemplars with detailed reasoning explanations (i.e., rationales [13, 42, 60, 91]) that lead to the final answer. However, manually curating such demonstrations is costly, as it requires extensive efforts and domain knowledge [98]. To mitigate this limitation, automatic chain-of-thought prompting (Auto-CoT) [96] is introduced to automatically select instances from the corpus coupled with zero-shot CoT [39], where the rationales are generated by LMs. Furthermore, it has been shown that CoT reasoning can be elicited even without explicit prompting, particularly for instruction-tuned LMs [75]. Another related work shows rationales generated by small models can help large models reason better [43]. However, none of them are designed for RAG, and how to leverage the instruction-following abilities of LMs for explicit denoising in the context of RAG is still underexplored.
@@ 5 Conclusion
In this work, we mainly conduct experiments on question answering-type tasks, and it remains unclear how our method may generalize to other scenarios (e.g., open-ended generation or multi-turn conversations that are also knowledge-intensive). Moreover, despite following the standard evaluation metrics, both accuracy and exact match are biased and cannot perfectly reflect the quality of the model’s generations. For instance, such metrics heavily rely on string matching, which assesses correctness at the lexical level rather than the semantic level, thereby failing to recognize different phrasings that convey identical meanings. The evaluation results also suffer from length bias, as longer generations tend to achieve higher accuracy. Although our method is not greatly affected by the aforementioned evaluation limitation, we acknowledge the necessity to explore more comprehensive metrics that can accurately evaluate model generations in RAG, such as using LLMs as judges [87]. Furthermore, we use a fixed number of retrieved documents under each setting as in line with prior studies. It will be interesting to evaluate the model performance under long-context settings with a dynamic or extremely large number of retrieved documents.
@@
Future research directions include exploring more advanced techniques for generating high-quality rationales, such as incorporating domain-specific knowledge or leveraging multi-task learning to enable better generalization across various tasks. For instance, although the consistency ratio between synthetic rationales and ground-truth answers on training samples with at least one relevant document achieves 98%, the overall consistency ratio on all training samples is only 89%. This is because for some samples, none of the retrieved documents is relevant to the question, which significantly compromises the quality of the generated rationales. Therefore, it will be interesting to fully explore the potential of our method by incorporating additional designs such as a filtering mechanism, which we leave as future work. Additionally, investigating the scalability of InstructRAG to larger datasets and models could further demonstrate its generalizability and robustness. Finally, integrating InstructRAG with other advanced retrieval techniques, such as active retrieval, could potentially lead to even better performance on knowledge-intensive tasks.

### v2

# InstructRAG: Instructing Retrieval-Augmented Generation via Self-Synthesized Rationales

@@ 1 Introduction
In light of this, retrieval-augmented generation (RAG) [3, 23, 29, 42, 52] has been introduced to enhance the generation accuracy of LMs in knowledge-intensive tasks by leveraging the most up-to-date information and specialized knowledge from external sources [40, 83, 102, 117]. However, the retrieved contents are typically mixed with irrelevant or even erroneous information due to the absence of perfect retrieval solutions [27, 39, 43, 44, 74] and the presence of noisy data in the retrieval corpus [28, 53, 105], posing a long-standing challenge to almost all RAG systems. Typically, vanilla RAG approaches address this issue implicitly by training LMs to directly predict correct answers despite potentially noisy inputs. Such latent processes are not only difficult to interpret and verify but also vulnerable to higher noise ratios, especially when the number of retrieved documents is large [11, 13, 56, 91]. On the other hand, obtaining high-quality explicit denoising supervision often requires substantial human efforts, which is time-consuming and costly.
@@ 2 Our Method: InstructRAG
As detailed in Algorithm 1, during the entire process, InstructRAG does not require any additional supervisions beyond standard RAG methods. By default, we instantiate both and with the same off-the-shelf instruction-tuned model (i.e., meta-llama/Meta-Llama-3-8B-Instruct), making InstructRAG a fully self-synthesis method. We also experiment with different instantiations of and and conduct ablation study in both training-free and trainable settings (§ 3.3). For simplicity, we use placeholders to represent omitted instructions in the prompts presented in this section, while the full list of complete prompt templates is provided in Appendix C.
@@ 2.1 Problem Setting
We adopt the standard RAG setting where the LM has access to annotated datasets of downstream tasks (e.g., question-answering task ), and an external knowledge base with the off-the-shelf retriever for retrieval. Different from previous works [3, 105] which leverage additional supervisions from GPT-3 [6] or GPT-4 [1], we assume the model has strictly limited access to the above two information sources. Given a question , the retriever returns a set of potentially noisy documents from the external knowledge base. The model is then tasked to predict the correct answer to the given question based on and its own parametric knowledge, denoted as .
@@ 2.2 Rationale Generation via Instruction-Following
Recent studies [51, 63, 67] have made encouraging progress in aligning LMs with human preferences and intentions, enabling the synthesis of high-quality data that closely follows user instructions [98]. Inspired by these advances, we propose to leverage the LM’s strong instruction-following ability to generate explicit denoising responses (i.e., rationales) for RAG. As shown in Table 1, given a QA pair and a set of retrieved documents , we prompt an off-the-shelf LM (as the rationale generator) with denoising instructions to produce the corresponding rationale that distinguishes useful documents from noisy ones and explains how the contexts lead to the ground-truth answer . To ensure the synthetic rationales are aligned with the ground-truth answers, we use a simple substring match to assess their consistency. The consistency ratio on training samples with at least one relevant document containing the ground-truth answer is 98% on average across five benchmarks, supporting the reliability of synthetic rationales as a sanity check. This allows us to effectively augment the standard dataset with self-synthesized denoising rationales solely by instructing the LM, without any additional supervision.
We also validate the necessity of using an LM-based generator (i.e., ) to create the rationales instead of employing simple heuristics — without the generator, rationales can be created in a template-based manner (Table 6), by roughly identifying relevant retrieved documents through simple substring-matching with the ground-truth answer. However, as demonstrated in our ablation study, this approach suffers from semantically inaccurate matching of relevant documents, leading to significant performance degradation. Another advantage of the LM-based generator is that it can produce high-quality rationales even without referring to the ground-truth answer, which only results in a minor performance drop. More detailed analyses on rationale generation design can be found in our ablation study (§ 3.3).
@@ 3.1 Experimental Setting
RAG tasks and evaluation metrics. We extensively validate the effectiveness of InstructRAG on five knowledge-intensive benchmarks, including PopQA [61], TriviaQA [38], Natural Questions [47], ASQA [78], and 2WikiMultiHopQA [24]. We use Wikipedia corpus as the retrieval source, and test our method with both sparse and dense off-the-shelf retrievers, including BM25 [71], DPR [39], GTR [65] and Contriver [27]. The retrieval quality is measured by Recall@, indicating whether the retrieved documents contain the correct answer. Table 2 shows the detailed dataset statistics. Following standard evaluation settings [3], we adopt the official metric of correctness (str-em), citation precision (pre) and recall (rec) for ASQA [18], and use accuracy for the other tasks, which measures whether the ground-truth answers are included in the model generations [61, 73]. Additionally, we also adopt LLM-as-a-judge for further evaluation (§ 3.4), as the above standard metrics are subject to the limitations of pattern-matching, which cannot accurately handle semantic equivalence.
Baselines. We compare our method with a wide range of RAG baselines under both training-free and trainable settings. Given that state-of-the-art LMs have incorporated a large amount of world-knowledge during the pre-training stage, we also report the performance of a non-retrieval baseline (namely, vanilla zero-shot prompting) for reference. Specifically, the training-free RAG baselines includes: (1) in-context retrieval-augmented language modeling (RALM) [70], a prompting method that extends the non-retrieval baseline by presenting the model with retrieved documents; (2) few-shot demonstration with instruction, an ICL method using ground-truth question-answer pairs sampled from the training set as demonstration exemplars.
The trainable RAG baselines include: (1) vanilla supervised fine-tuning (SFT), a supervised method with the training objective of maximizing the data likelihood of ground-truth answer given potentially noisy input; (2) RetRobust [105], which fine-tunes the RAG model on a mixture of relevant and irrelevant contexts to make it robust to irrelevant contexts; (3) Self-RAG [3], a strong trainable baseline, focusing on adaptive retrieval controlled by special reflection tokens. Both RetRobust and Self-RAG were originally built on Llama-2 [82] with additional supervisions. For example, RetRobust augments the training data for multi-hop reasoning tasks (e.g., 2WikiMultiHopQA) by prompting GPT-3 to decompose the original query and generate intermediate subqueries, and Self-RAG requires GPT-4 to generate additional reflective tokens to augment training samples.
For a fair comparison, we re-implement the two methods on Llama-2 and/or Llama-2 with augmented training data released by their authors, and report their performance as the higher one between the original scores and our reproduced results. As our method adopts instruction-tuned Llama-3 as the backbone model, we also train RetRobust and Self-RAG with Llama-3-Instruct and optimize their performance through extensive hyper-parameters search. More details on implementation, including training, inference, and prompt design are available in Appendix A and Appendix C.
@@ 3.2 Main Result
RAG without training. The second block shows the comparison among training-free RAG methods. In-context RALM and few-shot demonstration with instruction methods generally achieve higher performance than the non-retrieval baseline, highlighting the importance of retrieval for knowledge-intensive tasks. Encouragingly, our InstructRAG-ICL consistently outperforms all training-free baselines across various metrics, confirming the effectiveness of self-synthesized denoising rationales. Moreover, the boost from 8B to 70B model also demonstrates that InstructRAG-ICL scales effectively with larger backbone models, validating the generalizability of our method.
RAG with training. As present in the bottom block of Table 3, our InstructRAG-FT not only surpasses all non-retrieval and training-free baselines across all five benchmarks, but also significantly outperforms trainable RAG baselines on almost every metric. The only exception is in the ASQA task, where our method slightly underperforms Self-RAG in terms of citation (i.e., pre and rec). This is because our work primarily focuses on explicit denoising for RAG to improve the correctness of generations, which is measured by em. Despite not being explicitly optimized for citation metrics, our method still achieves competitive citation performance, significantly enhancing both generation accuracy and trustworthiness. Note that RetRobust achieves competitive performance on 2WikiMultiHopQA, which involves multi-hop reasoning. We attribute this to the additional training supervision provided by GPT-3, which enables the model to explicitly generate intermediate sub-queries and sub-answers. Another interesting finding is that Self-RAG consistently exhibits inferior performance compared to vanilla SFT, and even underperforms the training-free in-context RALM baseline across all benchmarks. We speculate the reason might be that these RAG tasks favor more domain-specific knowledge than general knowledge. However, it is challenging for Self-RAG to directly leverage in-domain features from existing training data as it requires GPT-4 to generate reflection tokens on these benchmarks, which is not available in our problem (§ 2.1).
3.3 Ablation Study
@@ 3.4 Analysis
Furthermore, we also study the generalizability of InstructRAG to a non-QA knowledge-intensive task such as code generation. As presented in Table 5(a), we directly apply InstructRAG-FT trained on the QA task (PopQA) to solve the unseen code generation task (HumanEval [12]), following the CodeRAG-Bench setup [87]. We evaluate the code generation performance using the standard pass@k metric and compare our method with the off-the-shelf Llama-3-8B-Instruct as the baseline. It can be observed that our method consistently achieves better generalization performance in the unseen code generation task in both non-retrieval and RAG settings. This finding aligns with our observation that InstructRAG trained on QA tasks tends to generate more text-based comments that articulate the design of coding solutions compared to the off-the-shelf Llama-3-8B-Instruct, thereby leading to more accurate code generation.
Evaluation with LLM-as-a-judge. Despite being standard evaluation metrics for question-answering, accuracy or exact match are known to be imperfect [13] as they mainly rely on pattern-matching to judge whether the predicted answer aligns with the ground-truth answer. Such metrics cannot handle cases where the predicted answer and ground-truth answer are synonyms (e.g., “Donald Trump” vs “Donald J. Trump” cannot be correctly recognized as a match), leading to biased evaluation results.
Therefore, we adopt the LLM-as-a-judge [7, 115] approach to evaluate whether the model response aligns with the ground-truth answer using GPT-4o [66], which allows the judge to consider semantic equivalence and is expected to yield a more fair evaluation. As shown in Table 5(b), we evaluate our method and baseline models on the open-domain Natural Questions benchmark in both training-free and trainable RAG settings. Compared to pattern-matching based metrics, LLM-as-a-judge generally leads to higher evaluation results, mostly due to its capability to accurately match semantically equivalent phrasings. Notably, our method consistently outperforms baselines under both pattern-matching based and LLM-based evaluation metrics, further validating the effectiveness of InstructRAG.
| Method | pass@1 | pass@10 |
| Without Retrieval | ||
| Llama-3-8B-Instruct | 58.5 | 64.6 |
| InstructRAG-FT | 60.4 | 65.2 |
| With Retrieval | ||
| Llama-3-8B-Instruct | 59.8 | 69.5 |
| InstructRAG-FT | 64.6 | 71.3 |
| Method | Pattern-based | LLM-based |
| RAG w/o Training | ||
| In-Context RALM | 56.8 | 64.5 |
| InstructRAG-ICL | 62.1 | 67.6 |
| RAG w/ Training | ||
| Vanilla SFT | 56.6 | 65.1 |
| InstructRAG-FT | 65.7 | 69.7 |
@@ 4.1 Retrieval-augmented Generation
Retrieval-augmented generation (RAG) is a widely adopted approach to enhance large language models (LLMs) with external knowledge [2, 4, 5, 19, 23, 42, 52, 70, 74], demonstrating promising potential to reduce hallucinations and enhance the generation accuracy of LLMs across various real-world applications [9, 37, 55, 58, 77, 80, 117]. Recently, a growing research effort has been devoted to enhancing RAG from various aspects, such as improving decoding efficiency [64, 36, 57], exploring long-context retrieval [96, 104], compressing prompts [32, 95], and addressing practical concerns such as adversarial retrieval [92, 116, 118] and privacy leakage [26, 108]. Despite their advantages, these RAG systems inevitably suffer from irrelevant information introduced by imperfect retrievers or noisy retrieval corpora. However, most existing works typically address this issue by improving the retrieval quality and reducing noise exposure to the model [22, 33, 72, 86, 100, 102, 109, 111]. Notable methods include adaptive retrieval [3, 35, 103] and query rewriting [8, 62]. In contrast, our work focuses on an orthogonal direction of developing explicit denoising methods for RAG, thereby enhancing the model’s noise robustness and generation accuracy, even in highly noisy contexts.
@@ 4.2 Eliciting Reasoning in Large Language Models
Recent studies have extensively explored the reasoning capability of LMs, but typically not in the context of RAG where potentially noisy retrieved contents may mislead the reasoning if not properly addressed. Chain-of-thought (CoT) prompting [89] is an effective method to elicit step-by-step reasoning from LMs by showing exemplars with detailed explanations (i.e., rationales [16, 49, 69, 107]) that lead to the final answer. However, such works often requires manually crafted demonstrations [84, 97], which is costly and requires extensive efforts and domain knowledge [114]. To mitigate this limitation, automatic chain-of-thought prompting (Auto-CoT) [112] is introduced to automatically select instances from the corpus coupled with zero-shot CoT [46], where the rationales are generated by LMs. Furthermore, it has been shown that CoT reasoning can be elicited even without explicit prompting, particularly for instruction-tuned LMs [85]. Another related work shows rationales generated by small models can help large models reason better [50]. Although rationalization has been extensively investigated in many NLP tasks [10, 20, 68, 90], none of them are designed for RAG, and how to leverage the instruction-following abilities of LMs for explicit denoising in the context of RAG is still underexplored.
@@ 5 Conclusion
In this work, we mainly conduct experiments on question answering-type tasks, and it remains unclear how our method may generalize to other scenarios (e.g., open-ended generation). Moreover, despite being the standard evaluation metrics, both accuracy and exact match are biased and cannot perfectly reflect the quality of the model’s generations. For instance, such metrics heavily rely on string matching, which assesses correctness at the lexical level rather than the semantic level, thereby failing to recognize different phrasings that convey identical meanings. The evaluation results also suffer from length bias, as longer generations tend to achieve higher accuracy. Exploring more advanced metrics like using LLMs as judges would better evaluate RAG model generations [102]. Another potential limitation is that our model might be subject to sample bias in the training data. Incorporating bias mitigation methods [17, 41, 101] would be helpful for further improving our work.
@@
Future research directions include exploring more advanced techniques for generating high-quality rationales, such as incorporating domain-specific knowledge or leveraging multi-task learning to enable better generalization across various tasks. For instance, although the consistency ratio between synthetic rationales and ground-truth answers on training samples with at least one relevant document achieves 98%, the overall consistency ratio on all training samples is only 89%. This is because for some samples, none of the retrieved documents is relevant to the question, which significantly compromises the quality of the generated rationales. Therefore, it will be interesting to fully explore the potential of our method by incorporating additional designs such as a filtering mechanism, which we leave as future work. Additionally, investigating the scalability of our method to larger datasets and models could further demonstrate its generalizability and robustness. It will also be interesting to evaluate the model performance under long-context settings with a dynamic or extremely large number of retrieved documents. Finally, integrating our method with other advanced retrieval techniques, such as active retrieval, could potentially lead to even better performance on knowledge-intensive tasks.